The Westwood Warriors basketball team wants to score more points. To get better at scoring points the team is trying to improve its offensive strategies. Some opponents primarily use a zone defense, while others primarily use a man-to-man defense. When the Warriors play against teams that use a zone defense they score an average of 67 points per game. When they play against teams that use a man-to-man defense they score an average of 62 points per game. The standard deviations are 8 and 5 respectively. Since the Warriors started using their improved offensive strategies, they have played two games with the following results:
Against the McNeil Mavericks Against the Round Rock Dragons
Maverick defense: zone Dragon defense: man-to-man
Warrior points: 77 Warrior points: 71
Which offensive strategy -- learning to play against the zone defense or learning to play against the man-to-man defense – appears to have improved the most? (Of course, there are many factors involved in playing basketball. For your answer, assume the difference is due to the new offensive strategies.)
Explain Completely.
Answer #1

ANSWER: Aganist McNeil Mavericks. M = 6+ r-8 X: 17 Z = X-94 Solo) 47-67 B 1.25 P-vaklee 0.8943 Against the Round ROCK Doagons